# Evaluating Grounded Reasoning by Code-Assisted LLMs for Mathematics

### v1

@@ Acknowledgments
The authors would like to thank Vagrant Gautam for their mentorship and useful feedback. We are also grateful to Marius Mosbach and Ellie Pavlick for feedback during the early stages of this work. We also thank Miaoran Zhang for the useful discussions and Tural Mammadov for help in the experimental infrastructure. The authors received funding from the DFG (German Research Foundation) under project 232722074, SFB 1102.
@@ C.2 Examples of incorrectly classified programs
def smallest_page_number(): # The least common multiple (LCM) of 4 and 13 lcm = 4 * 13 # Since 4 and 13 are co-prime return lcm result = smallest_page_number() True label: “No Logic”
@@ D.1 Most common bugs in the generated programs on MATH500 dataset
D.2 Analysis of programs per math subject
Figure 9 illustrates the distribution of taxonomy classes per math subject from the MATH500 dataset.
(a) GPT4o-mini
(b) Qwen2.5
(c) StarCoder2
Figure 9: Distribution of taxonomy classes counts per math subject from the MATH500 dataset for evaluated models. Classes: ’FS Imp’: From-Scratch Implementation, ’BF’: Brute-Force ’Disorg’: Disorganized.

### v2

@@ Acknowledgments
The authors would like to thank Vagrant Gautam for their mentorship and useful feedback. We are also grateful to Marius Mosbach and Ellie Pavlick for feedback during the early stages of this work, to Miaoran Zhang for the valuable discussions, and to Tural Mammadov for help with the experimental infrastructure. We also thank anonymous reviewers for feedback. The authors received funding from the DFG (German Research Foundation) under project 232722074, SFB 1102.
@@ C.2 Examples of incorrectly classified programs
def solution(): cards_per_pack = 20 envelopes_per_pack = 17 # least common multiple of 20 and 17 is 340 lcm = 340 result = lcm return result solution() True label: “No Logic”